One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by checkley77(12844)

X=Y+12
Y+2X=39
Y+2(Y+12)=39
Y+2Y+24=39
3Y=39-24
3Y=15
Y=15
Y=5 ANS.
X=5+12=17 ANS.
PROOF:
5+2*17=39
5+34=39
39=39